Prev Next

SW&IT Meeting Minutes 2019-09-09

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20190826
LR20190826 - Bec, Matthieu 72.0
LR20190826 - Holmberg, Carl 80.0
LR20190826 - Colson, Andrew 64.0
LR20190826 - Edmeades, David 70.0
LR20190826 - Sola, Igor 80.0
LR20190826 - Kubanek, Petr 82.9
LR20190826 - Hooper, Stephen 72.0
LR20190826 - Zhang, Yang 72.0
LR20190826 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Open topics

  • CI/CD (Gitlab, GitLab license, wait for GitHub Actions runner, keep on Jenkins, move to GitLab, find something else) [PK]
    • LBTO not eligible to gitlab/edu
    • github/CI in beta, release Nov 13? mark our calendars
  • Wiki woes (ctd.)
    • if anything, it's getting worse.
    • reinstall/upgrade OS, more probes?
  • IT issues, email limitations [MB]
  • UITS coloc update
  • IT support while SH on leave
  • ObsWsUp, x2go/zoom - MATE?

Team reports

MB

  • OC, WPC, Management
  • PP2020 discussion
  • Steward/GMT M1 visit UA-tech park
  • DMS
    • work on M1 surfaces
    • F30 fixes, mountain fixes
    • new Alta stream, mongodb adapter
  • Alta short timescale predictions meeting
  • safety training
next
  • review YZ's AO-cravat POI
  • phantom calls follow-up, who?
  • DMS
    • Alta
    • m1surf
    • short-link
  • Argos issue doc.

YZ

  • AO-Cravat:
    • Put together a draft version of the POI (in Google doc, now shared with relevant people).
    • Met with Steve Ertel to discuss his Strehl program.
    • Next is to work on refactoring Steve's code and create Github entry.
    • Also to investigate Chart.js plotting (or an alternative)
    • Keep editing POI when needed.
  • MODS/LBC:
    • Worked on IT7517 (MODS1 WFS camera problem) with Petr.
    • Created an initial Python program to ingest MODS ISIS telemetry.
    • Created relevant MongoDB for MODS telemetry.
    • Forked and modified the web-dms code, initial testing successful.
    • Next is to complete the whole setup for MODS, and then do the same for LBC housekeeping telemetry.
    • Discuss LBC FPIA/WRC code performance with Olga.
    • Several IT issue tracks to follow up with Olga.
  • Completed all safety training courses!

XZ

Last week:
  • Support the ASM flat issue
  • Support LBTI SXWFS configuration
  • AO telemetry requirements(scope)
  • Backup the AdSec Recs
  • AdSec current saturation control
  • Install git on flao-dxwfs
  • Fix some 4D issues
This week:
  • AO support
  • Support LBTI SXWFS configuration
  • Safety training

DE

  • Wiki daily reboots
    • Unsure what happened Saturday
  • 507 display plan?
  • Colocation racks ordered
  • Vault backup email notifications sorted; don't play games with the domain or Google might decide to blackhole them
  • Progress with VMWare, have new people working on getting me a quote
  • Wiki virtual hardware tweaks
  • DMS mounts/systemd/Github
Next
  • Nail down VMWare
  • Monolithic PR for telemetry mounts
  • Minor changes to DMS/mountain ks for mounts
  • Wiki monitoring

SH

  • browser extension for bookmarks.
  • kickstart of remote observing VM
  • Fix docker mount for Petr
  • fix xgterm for John
  • re-jigger radar image
  • RPMs for mount points
  • support calls for a few things
next:
  • add users from ssh.lbto.org to remote observing VM
  • move ssh.lbto.org home directories to share
  • VNC on rm507 computers
  • observer reset scripts
  • RPM packaging of bookmarks, obs reset script
  • KS DMS

PK

  • TCS 2019A
  • Qt for MCU (microcontrolers) webinar
  • FITSpp: keyword manipulation (delete, update,..), uint64_t fix
  • Learned a lot about various CamLink fibers and their points of failure (IT7834, IT7517)
  • Small TCS patches after 2019A (gcsclient, missed GCS IT7643 addition, fixed ENVGUI colors, IT7845)
  • Toured Carl a bit around remote room during night operations

CH

  • Last week:
    • Witnessed a couple hours of night ops.
    • Broke out vibration data into individual fields in iMonnit script.
    • Attempted to add temperature to dms backend, but suffered from not keeping up with MB commits.
  • This week:
    • Merge MB dms changes.
    • Finish dms-monnit backend additions.

AC

  • IT7612 - updated FacSum with new radar image (Thanks SH!)
  • IT7831 - reinstalled tcsweb RPM and restarted facsum.sh on mountain middle monitor. Displayed as expected before power outage but then went grey screen with "label_34'"message.
  • Overhaul FacSum UI - very difficult to resize and place widgets due to poor design. Re-organizing layouts.

next:
  • IT7848 - re-enable vncserver on overhead displays
  • continue and test FacSum redesign improvements
  • AGW/OAC ALH - check alarm for UMAC

IS

Last week
  • Luci
    • Finish command line option for taking grating calibration for Dave T.
    • Looking at issues with RTD mask tab, particularly the target error calculation. (ongoing)
    • Issues with firing the Luci scripts. Don't understand the issues. Don't know if they are new or operator errors? Created an lbto account in lucix for users to run luci client software, waiting to see what the new complaints are.
  • OT
    • Providing a table with filter capabilities will all the proposals Ids, pi names and emails
    • Fixing some of the MODS scripting issues
    • other bugs

This week:
  • OT
    • Keep working on MODS scripting issues
    • Finishing proposal id work
    • other bugs
    • hopefully have the release ready by Friday
  • LUCI
    • look at RTD MOD tab error calculation
Topic revision: r16 - 09 Sep 2019, AndrewColson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback